# Goal-Oriented Communications for Real-time Inference with Two-Way Delay

### v1

@@ II-B Inference Error
Figure 2: A sample evolution for two consecutive transmission periods.
Let $d(t)$ represent the length of the most recently delivered packet by time slot $t$, which can be written as follows:
@@
We focus on a class of widely used supervised learning algorithms known as Empirical Risk Minimization [42]. We assume that, for each packet length $l\in\{1,2,\ldots,B\}$, the receiver contains a trained neural network $\phi^{*}_{l}\big{(}\cdot\big{)}$ that takes the AoI $\Delta(t)\in\mathbb{Z}^{+}$ and the packet $X_{t-\Delta(t)}^{l}\in\mathcal{V}^{l}$ as inputs and produces an output $a\in\mathcal{A}$ to infer the real-time value of the target signal $Y_{t}\in\mathcal{Y}$. The performance is evaluated through a loss function $L\big{(}\cdot\big{)}:\mathcal{Y}\times\mathcal{A}\mapsto\mathbb{R}$. The incurred loss is $L(y,a)$ when the output $a$ is taken to predict $Y_{t}=y$. Given the AoI $\delta$ and packet length $l$, the training problem to obtain the neural network $\phi^{*}_{l}\big{(}\cdot,\cdot\big{)}$ is formulated as:
$\min_{\phi\big{(}\cdot,\cdot\big{)}\in\Lambda}\mathbb{E}_{Y,X^{l}\sim P_{\tilde{Y}_{0},\tilde{X}_{-\delta}^{l}}}\big{[}L\big{(}Y,\phi\big{(}\delta,X^{l}\big{)}\big{)}\big{]},$ (3)
where $\Lambda$ is the set of all mappings that the neural network can generate, and $P_{\tilde{Y}_{0},\tilde{X}_{-\delta}^{l}}$ represents the empirical distribution of the target signal $\tilde{Y}_{0}$ and the packet $\tilde{X}_{-\delta}^{l}$. The AoI $\delta$ is the time difference between the generation of $\tilde{Y}_{0}$ and $\tilde{X}_{-\delta}$.
These trained predictors are used to estimate the real-time value of the target signal $Y_{t}$ on the receiver side. We assume that the process $\{\big{(}Y_{t},V_{t}\big{)},t=0,1,\ldots\}$ is stationary, and that the processes $\{\big{(}Y_{t},V_{t}\big{)},t=0,1,\ldots\}$ and $\{\big{(}\Delta(t),d(t)\big{)},t=0,1,\ldots\}$ are independent. The average inference error at time slot $t$, given $\Delta(t)=\delta$ and $d(t)=l$, is expressed as:
@@
We are interested in loss functions $L\big{(}\cdot\big{)}$ for which the inference error (4) is bounded. Apart from this restriction, the loss function $L\big{(}\cdot\big{)}$ can be chosen based on the goal of the remote inference system. For example, a quadratic loss function $L_{2}\big{(}y,\hat{y}\big{)}=\lVert y-\hat{y}\rVert_{2}^{2}$ is used in neural network-based minimum mean-squared estimation, where the action $a=\hat{y}$ is an estimate of the target signal $Y_{t}=y$ and $\lVert y\rVert_{2}$ is the Euclidean norm. In softmax regression (i.e., neural network-based maximum likelihood classification), the action $a=\mathcal{Q}_{Y}$ is a distribution of $Y_{t}$, and the loss function $L_{\log}\big{(}y,\mathcal{Q}_{Y}\big{)}=-\log\mathcal{Q}_{Y}(y)$ is the negative log-likelihood function of the value $Y_{t}=y$.
In the subsequent sections, we solve a learning and communication co-design problem that aims to optimize the performance of a remote inference system. From equation (4), we know that the inference error is a function of both the AoI and packet length. For a given AoI $\delta$, the inference error is a non-increasing function of packet length [17]. However, for a given packet length $l$, the inference error does not always increase monotonically with the AoI [4, 5]. This observation contradicts the conventional assumption that fresher packets yield smaller inference errors. Theoretical and experimental explanations of this counter-intuitive observation was provided in [4, 5]. Therefore, we do not adopt the generate-at-will model [18, 19] in this paper.
II-C Structure of a Scheduling Policy
Upon the ACK reception time $A_{i}$, the scheduler determines a waiting time $\tau_{i+1}$ to specify the next submission time slot $S_{i+1}=A_{i}+\tau_{i+1}$. Then, a data packet $X_{S_{i+1}-b_{i+1}}^{l_{i+1}}$ is formed by selecting the packet length $l_{i+1}$ and buffer position $b_{i+1}$. A scheduling policy $\pi\in\Pi$ is defined as a tuple $\pi=\big{(}f,\ell,g\big{)}$, where $f=\big{(}b_{2},b_{3},\ldots\big{)}$ is the buffer position sequence, $\ell=\big{(}l_{2},l_{3},\ldots\big{)}$ is the packet length sequence, $g=\big{(}\tau_{2},\tau_{3},\ldots\big{)}$ is the waiting time sequence, and $\Pi$ denotes the set of all causal policies $\pi$. We assume that the scheduling policy does not employ the knowledge of the observed process $\{\big{(}Y_{t},V_{t}\big{)},t=0,1,\ldots\}$. This assumption implies that the processes $\{\big{(}Y_{t},V_{t}\big{)},t=0,1,\ldots\}$ and $\{\big{(}\Delta(t),d(t)\big{)},t=0,1,\ldots\}$ are independent. The initial conditions of the system are assumed to be $S_{1}=0$, $l_{1}=1$, $b_{1}=0$, $c_{1}\in\{1,2,\ldots,N\}$, and $\Delta(0)$ is a finite constant.
@@ III Learning and Communications Co-design: Time-invariant Packet Length Selection
While designing goal-oriented communication protocols for remote inference, selecting a constant packet length is a simple and often effective choice. If packet lengths vary dynamically from one packet to the next, the receiver needs to reconfigure the estimator over time to accommodate the changing input dimensions, which adds complexity. In this section, we consider systems where this complexity is avoided by choosing a time-invariant packet length sequence. Systems that can vary the packet length over time are discussed in Section IV.
@@ III-A Co-design Problem Formulation
The learning and communications co-design problem with time-invariant packet length selection can be formulated as a two-layer nested optimization problem:
@@
The inner optimization problem (6) can be cast as an infinite-horizon average-cost Semi-Markov Decision Process (SMDP). While dynamic programming algorithms, namely policy iteration and value iteration, are typically employed to solve such problems [20, 21], we are able derive a closed-form solution that achieves significantly lower computational complexity than dynamic programming. Moreover, the outer optimization problem (7) is a straightforward search over the integer values $\{1,2,\ldots,B\}$.
Unlike the formulation in [17], this problem formulation accounts for varying delay conditions between the transmitter and the receiver and does not assume immediate feedback.
@@ III-B Optimal Solution to (6)-(7)
We define each ACK reception time $A_{i}$ as a decision time for the SMDP (6). The system state at time slot $A_{i}$ is represented by a tuple $\big{(}\delta,c_{i}\big{)}$, where $\delta=\Delta(A_{i})$ and $c_{i}$ is the delay state in the $i$-th epoch. The actions are the waiting time $\tau_{i+1}$ and the buffer position $b_{i+1}$. A detailed description of the SMDP can be found in Appendix A. The Bellman optimality equation of this SMDP can be formulated as:
@@
where $h_{1}(\delta,c_{i})$ is the relative value function. We need to solve the Bellman optimality equation (III-B) to find an optimal solution to (6). This involves determining the optimal buffer position $b_{i+1}^{*}$ and waiting time $\tau_{i+1}^{*}$ for each state $\big{(}\delta,c_{i}\big{)}$. By exploiting the structural properties of the Bellman optimality equation (III-B), we are able to obtain a closed-form solution to (6), as asserted in the following theorem.
@@
for all $\delta\in\{1,2,\ldots\}$, $d,l\in\{1,2,\cdot,B\}$, and $c\in\{1,2,\ldots,N\}$.
@@
$\mathbb{E}\Bigg{[}\sum_{t=A_{i}(\beta)}^{A_{i+1}(\beta)-1}\varepsilon\big{(}\Delta(t)\big{)}\Bigg{]}-\beta\mathbb{E}\big{[}A_{i+1}(\beta)-A_{i}(\beta)\big{]}=0,$ (12)
and it is exactly the optimum value $\varepsilon_{l,\text{opt}}$ of problem (6).
@@
Step 1: We obtain two key results by using the Bellman optimality equation (III-B):
  • The optimal waiting time $\tau_{i+1}^{*}$ is determined by the threshold rule $\tau_{l}\big{(}\delta,c_{i}\big{)}$ given by equation (10) for all $i$.
@@
where $\vartheta_{i+1}=F_{i+1}+\tau_{l}\big{(}b_{i+1}+T_{i+1}(l)+F_{i+1},C_{i+1}\big{)}+T_{i+2}(l)$, and $h_{1}^{\prime}\big{(}c_{i}\big{)}$ is the relative value function.
The Bellman optimality equation (III-B) is decomposable and can be solved as a per-decision-epoch optimization problem, allowing any $b_{i+1}^{*}$ in the optimal buffer position sequence ${f^{*}=\big{(}b_{2}^{*},b_{3}^{*},\ldots\big{)}}$ to be expressed as shown in (1).
Step 3: We determine that the optimal value $\varepsilon_{l,\text{opt}}$ for the inner optimization problem (6) is the unique root of (12).
@@
The optimal scheduling policy for the inner optimization problem (6), outlined in Theorem 1, is well-structured. Each optimal waiting time $\tau_{i+1}^{*}$ is determined by an index-based threshold rule $\tau_{l}\big{(}\delta,c_{i}\big{)}$. The index function $\gamma\big{(}\cdot\big{)}$, provided in (9), can be readily calculated for any state $\big{(}\delta,c_{i}\big{)}$, using the conditional distribution of $T_{i+1}(l)$ given $C_{i}=c_{i}$.
Furthermore, according to equation (1), $b_{i+1}^{*}$ is designed to minimize the relative inference error in the interval between $D_{i+1}$ and $D_{i+2}$, given that the waiting time is determined by the optimal rule in (10). Also, each optimal buffer position $b_{i+1}^{*}$ depends solely on the delay state $c_{i}$ in the $i$-th epoch, and is independent of $\delta=\Delta(A_{i})$. This result can be better understood by deeply analyzing the AoI evolution. At each delivery time slot $D_{i+1}$, the AoI resets to a realization of $b_{i+1}^{*}+T_{i+1}(l)$. Due to these regular resets, the AoI evolution between consecutive delivery slots $D_{i+1}$ and $D_{i+2}$ is independent of the AoI evolution outside this interval. Moreover, $b_{i+1}^{*}$ is a parameter to control the AoI evolution between $D_{i+1}$ and $D_{i+2}$.
Finally, the optimum value $\varepsilon_{l,\text{opt}}$, which is the threshold in the waiting time rule (10) and necessary for solving (1), is the unique root of equation (12). This equation can be efficiently solved using low-complexity algorithms [25, Algorithms 1-3], such as bisection search.
The optimal policy in Theorem 1 has two key distinctions compared to the policy established in [17, Theorem 1] for remote inference under time-invariant packet length selection, IID transmission delays, and immediate feedback. First, although the threshold rule $\tau_{l}\big{(}\delta,c_{i}\big{)}$ in (10) closely resembles the waiting time rule derived in [17, Theorem 1], it is not solely dependent on the AoI but also incorporates the delay state. For the problem formulation (6)-(7), which involves a finite number of delay states with non-zero feedback delay, the existence of such a waiting time rule was not evident and required additional technical efforts for its derivation.
Second, in [17, Theorem 1], the optimal buffer position sequence is time-invariant due to the assumption of IID transmission delays. However, when the delay state changes with memory, achieving optimal performance requires a buffer position sequence that can adapt to variations in delay statistics. Theorem 1 unveils that the optimal buffer position sequence is determined by solving $N$ per-sample optimization problems, each related to a distinct delay state, without requiring pre-computation of the relative value function $h_{1}\big{(}\cdot\big{)}$.
We can solve the inner optimization problem (6) for any given packet length $l$ by using Theorem 1. Then, the outer optimization problem (7) is just a straightforward search over the integer values $\{1,2,\ldots,B\}$, and $\varepsilon_{\text{opt}}^{\text{inv.}}=\varepsilon_{l^{*},\text{opt}}$.
@@ IV Learning and Communications Co-design: Time-Variable Packet Length Selection
In contrast to the previous case, where high computational complexity is avoided, here we aim to achieve optimal performance without any restrictions on the packet length sequence $\ell=\big{(}l_{2},l_{3},\ldots\big{)}$. This unconstrained approach allows us to maximize the system’s potential by dynamically adjusting the packet length $l_{i}$ over time. Therefore, this case represents the most general solution.

### v3

@@ II-B Inference Error as a Function of AoI and Packet Length
Let $d(t)$ represent the sequence length of the most recently delivered packet by time slot $t$, which is given by
@@
We assume that the predictor employs a supervised learning algorithm based on Empirical Risk Minimization [42]. For each packet length $l\in\{1,2,\ldots,B\}$, the receiver contains a pre-trained neural network $\phi^{*}_{l}:\mathbb{Z}^{+}\times\mathcal{V}^{l}\mapsto\mathcal{A}$, which takes the AoI $\Delta(t)$ and the packet $X_{t-\Delta(t)}^{l}$ as inputs and produces an action $a$ to infer the real-time value of the target signal $Y_{t}$. The performance is evaluated through a loss function ${L:\mathcal{Y}\times\mathcal{A}\mapsto\mathbb{R}}$. The incurred loss is $L(y,a)$ when the action $a$ is taken to predict $Y_{t}=y$. Given the AoI $\delta$ and packet length $l$, the training problem to obtain the neural network $\phi^{*}_{l}$ is formulated as:
$\phi^{*}_{l}=\operatorname*{arg\,min}_{\phi\in\Lambda}\mathbb{E}_{Y,X^{l}\sim P_{\tilde{Y}_{0},\tilde{X}_{-\delta}^{l}}}\big{[}L\big{(}Y,\phi\big{(}\delta,X^{l}\big{)}\big{)}\big{]},$ (3)
where $\Lambda$ is the set of all mappings that the neural network can generate, and $P_{\tilde{Y}_{0},\tilde{X}_{-\delta}^{l}}$ represents the empirical distribution of the target signal $\tilde{Y}_{0}$ and the packet $\tilde{X}_{-\delta}^{l}$ in the training dataset. The AoI $\delta$ denotes the time difference between the generation of $\tilde{Y}_{0}$ and $\tilde{X}_{-\delta}^{l}$.
These trained neural networks are used to predict the real-time value of the target signal $Y_{t}$ on the receiver side. We assume that the process ${\{\big{(}Y_{t},V_{t}\big{)},t=0,1,\ldots\}}$ is stationary, and that the processes ${\{\big{(}Y_{t},V_{t}\big{)},t=0,1,\ldots\}}$ and ${\{\big{(}\Delta(t),d(t)\big{)},t=0,1,\ldots\}}$ are independent. The average inference error at time slot $t$, given $\Delta(t)=\delta$ and $d(t)=l$, is expressed as:
@@
The loss function $L$ can be chosen based on the goal of the remote inference application. For example, a quadratic loss function $L_{2}\big{(}y,\hat{y}\big{)}=\lVert y-\hat{y}\rVert_{2}^{2}$ is used in neural network-based minimum mean-squared estimation, where the action $a=\hat{y}$ is an estimate of the target signal $Y_{t}=y$ and $\lVert y\rVert_{2}$ is the Euclidean norm. In softmax regression (i.e., neural network-based maximum likelihood classification), the action $a=\mathcal{Q}_{Y}$ is a distribution of $Y_{t}$, and the loss function $L_{\log}\big{(}y,\mathcal{Q}_{Y}\big{)}=-\log\mathcal{Q}_{Y}(y)$ is the negative log-likelihood function of the value $Y_{t}=y$.
In the subsequent sections, we solve a learning and communication co-design problem that aims to optimize the performance of a remote inference system. From equation (4), we know that the inference error is a function of both the AoI and packet length. For a given AoI $\delta$, the inference error is a non-increasing function of packet length [17]. On the other hand, for a given packet length $l$, the inference error does not always increase monotonically with the AoI [5, 6]. When minimizing non-decreasing functions of AoI, fresher data packets are preferable, and the generate-at-will medium access model works well. However, to minimize inference error, which is not necessarily non-decreasing with respect to AoI, the selection-from-buffer model—allowing for the choice between fresh and older data packets—is more effective.
II-C Structure of the Scheduling Policy
Upon receiving the ACK at time slot $A_{i}$, the scheduler determines a waiting time $\tau_{i+1}$ to specify the next submission time slot as $S_{i+1}=A_{i}+\tau_{i+1}$. Then, a new data packet $X_{S_{i+1}-b_{i+1}}^{l_{i+1}}$ is formed with packet length $l_{i+1}$ and relative buffer position $b_{i+1}$. A scheduling policy $\pi\in\Pi$ is defined as a tuple $\pi=\big{(}f,\ell,g\big{)}$, where $f=\big{(}b_{2},b_{3},\ldots\big{)}$ is the relative buffer position sequence, $\ell=\big{(}l_{2},l_{3},\ldots\big{)}$ is the packet length sequence, $g=\big{(}\tau_{2},\tau_{3},\ldots\big{)}$ is the waiting time sequence. In addition, let $\Pi$ denote the set of all causal policies $\pi$. We assume that the scheduler does not utilize knowledge of the process $\{\big{(}Y_{t},V_{t}\big{)},t=0,1,\ldots\}$. This assumption implies that the processes $\{\big{(}Y_{t},V_{t}\big{)},t=0,1,\ldots\}$ and $\{\big{(}\Delta(t),d(t)\big{)},t=0,1,\ldots\}$ are mutually independent. The initial conditions of the system are assumed to be $S_{1}=0$, $l_{1}=1$, $b_{1}=0$, $c_{1}\in\{1,2,\ldots,N\}$, and $\Delta(0)$ is a finite constant.
@@ III Learning and Communications Co-design: Time-invariant Packet Length Selection
While designing goal-oriented communication protocols for remote inference, selecting a constant packet length is a simple and often effective choice. If packet lengths vary dynamically from one packet to the next, the receiver needs to reconfigure the predictor over time to accommodate the changing input dimensions, which adds complexity. This section considers systems where this complexity is avoided by choosing a time-invariant packet length sequence. Systems that can vary the packet length over time will be discussed later in Section IV.
@@ III-A Co-design Problem Formulation
The packet length affects both learning and communication. If it is selected based solely on learning performance, communication may suffer since transmission delays increase with packet length. Therefore, we solve a learning and communications co-design problem under time-invariant packet length selection, formulated as the following two-layer nested optimization problem.
@@
The inner problem (6) optimizes the relative buffer position sequence $f$ and waiting time sequence $g$ for a given packet length $l\in\{1,2,\ldots,B\}$. This problem can be cast as an infinite-horizon average-cost Semi-Markov Decision Process (SMDP). Although dynamic programming algorithms, such as policy iteration and value iteration, are typically used to solve such problems [23, 24], we are able to derive a closed-form solution that achieves significantly lower computational complexity than dynamic programming. Moreover, the outer problem (7) optimizes the packet length and involves a straightforward search over the integer values $\{1,2,\ldots,B\}$.
Problem (6)-(7) is a more general formulation than that considered in [17, Section IV-A], as it accounts for both (i) the two-way delay between the transmitter and receiver, and (ii) the Markovian, time-varying nature of the two-way delay distribution in multi-route communication networks.
@@ III-B Optimal Solution to (6)-(7)
Each time slot $A_{i}$ is a decision time for the SMDP (6). The system state at each $A_{i}$ is given by a tuple $\big{(}\Delta(A_{i}),C_{i}\big{)}$, where $\Delta(A_{i})$ represents the AoI at time-slot $A_{i}$, and $C_{i}$ indicates the delay state in the $i$-th packet transmission epoch. A realization of the system state is denoted by $\big{(}\delta,c_{i}\big{)}$. The actions determined at time slot $A_{i}$ are the waiting time $\tau_{i+1}$ and the buffer position $b_{i+1}$. A detailed description of this SMDP can be found in Appendix A. The Bellman optimality equation of the SMDP (6) is formulated as
@@
where $\delta\in\mathbb{Z}^{+}$, $c_{i}\in\{1,2,\ldots,N\}$, and $h_{1}(\delta,c_{i})$ is the relative value function. We need to solve the Bellman optimality equation (III-B) to find an optimal solution to (6). This involves determining the optimal buffer position $b_{i+1}^{*}$ and waiting time $\tau_{i+1}^{*}$ for each state $\big{(}\delta,c_{i}\big{)}$. By exploiting the structural properties of the Bellman optimality equation (III-B), we are able to obtain a closed-form solution to (6), as asserted in the following theorem.
@@
for all $\delta\in\mathbb{Z}^{+}$, $d,l\in\{1,2,\ldots,B\}$, and $c\in\{1,2,\ldots,N\}$.
@@
$\mathbb{E}\Bigg{[}\sum_{t=A_{i}(\beta)}^{A_{i+1}(\beta)-1}\varepsilon\big{(}\Delta(t),l\big{)}\Bigg{]}-\beta\mathbb{E}\big{[}A_{i+1}(\beta)-A_{i}(\beta)\big{]}=0,$ (12)
and $\beta=\varepsilon_{l,\text{opt}}$ is exactly the optimum value of problem (6). $A_{i}(\beta)$ in equation (12) represents the $i$-th ACK reception time slot when the scheduling policy is determined according to the rules specified by equations (10) and (1) for a given $\beta$.
@@
Step 1: We first establish two key results using the Bellman optimality equation (III-B):
  • The optimal waiting time $\tau_{i+1}^{*}$ is determined by the index-based threshold rule
    $\min\{k\geq 0:\gamma\big{(}\delta+k,l,l,c_{i}\big{)}\geq\varepsilon_{l,\text{opt}}\}$
    for all $i$. Here, $\gamma\big{(}\cdot\big{)}$ is the index function defined in (9).
@@
where $c_{i}\in\{1,2,\ldots,N\}$, $h_{1}^{\prime}\big{(}c_{i}\big{)}$ is the relative value function, and $\vartheta_{i+1}=F_{i+1}+\tau_{l}\big{(}b_{i+1}+T_{i+1}(l)+F_{i+1},C_{i+1}\big{)}+T_{i+2}(l)$.
Only the first term on the right-hand side of equation (III-B) depends on the action $b_{i+1}$. Thus, the Bellman optimality equation (III-B) is decomposable and can be solved as a per-decision-epoch optimization problem, allowing any $b_{i+1}^{*}$ in the optimal buffer position sequence ${f^{*}=\big{(}b_{2}^{*},b_{3}^{*},\ldots\big{)}}$ to be expressed as shown in equation (1).
Step 3: Finally, we show that the optimal value $\varepsilon_{l,\text{opt}}$ of problem (6) is the unique root of equation (12).
@@
The optimal scheduling policy for the inner optimization problem (6), as outlined in Theorem 1, is well-structured. Each optimal waiting time $\tau_{i+1}^{*}$ is determined by an index-based threshold rule $\tau_{l}\big{(}\delta,c_{i}\big{)}$. The index function $\gamma\big{(}\cdot\big{)}$, defined in equation (9), can be readily computed for any state $\big{(}\delta,c_{i}\big{)}$ using the conditional distribution of $T_{i+1}(l)$ given $C_{i}=c_{i}$. Although the index-based threshold rule $\tau_{l}\big{(}\delta,c_{i}\big{)}$ closely resembles the waiting time rule provided in [17, Theorem 1], it is not solely dependent on the AoI but also incorporates the delay state $C_{i}=c_{i}$. Because problem (6)-(7) involves non-zero feedback delay and time-varying delay distribution, the existence of such a simple waiting time rule was not evident and required additional technical efforts for its derivation.
Furthermore, according to equation (1), the optimal buffer position $b_{i+1}^{*}$ is designed by minimizing the relative inference error in the interval between $D_{i+1}$ and $D_{i+2}$, given that the waiting time is determined by the optimal rule in (10). Each $b_{i+1}^{*}$ depends solely on the delay state $C_{i}=c_{i}$ in the $i$-th epoch and is independent of the AoI state $\delta=\Delta(A_{i})$. For the case of i.i.d. transmission delay and immediate feedback, the optimal buffer position sequence is time-invariant, as shown in [17, Theorem 1]. However, when the feedback delay is non-zero and the delay state $C_{i}$ changes according to a Markov chain, achieving optimal performance requires that the optimal buffer position $b_{i+1}^{*}$ is a function of the delay state $C_{i}=c_{i}$, and this function is exactly given by (1).
Finally, the optimum value $\varepsilon_{l,\text{opt}}$, which is the threshold in the waiting time rule (10) and necessary for solving (1), is the unique root of equation (12). This equation can be efficiently solved using low-complexity algorithms [26, Algorithms 1-3], such as bisection search.
After solving the inner optimization problem (6) using Theorem 1, the outer optimization problem (7) is just a straightforward search over the integer values $\{1,2,\ldots,B\}$.
@@ IV Learning and Communications Co-design: Time-Variable Packet Length Selection
In contrast to the previous case of time-invariant packet length, this section aims to achieve optimal performance without imposing restrictions on the packet length sequence $\ell=\big{(}l_{2},l_{3},\ldots\big{)}$. This unconstrained approach allows us to maximize the system’s potential by dynamically adjusting the packet length $l_{i}$ over time. Therefore, this case represents the most general solution.